In a nuclear collision, a proton A is incident on a stationary neighboring proton B with a velocity of 3.00 x 10 m/s. After impact, A moves in the direction BC inclined at 600 with a velocity of 1.50 x 102 m/s. B then moves in the direction BD declined at 0. If BC makes an angle 600with the initial direction (AB) of A and BD makes an angle G with it ;
ii) Verify whether or not the collision is elastic.
1
Expert's answer
2020-11-05T10:12:27-0500


Let a proton "A" of mass "m" and velocity "\vec {u_A}" collides with another stationary proton "B" .

After collision the proton "A" moves with velocity "\vec {v_A}" at an angle "\theta" with the initial direction AB and the proton "B" moves with velocity "\vec{v_B}" at an angle "\phi" with AB as shown in the figure.

Conservation of linear momentum:

Conservation of linear momentum along x-axis:

"mu_A+m\times 0=mv_A\cos\theta+mv_B\cos\phi \\
\Rightarrow v_B\cos\phi =u_A-v_A\cos\theta \qquad ........(1)"

Conservation of linear momentum along y-axis:

"mv_A\sin\theta=mv_B\sin\phi \\
\Rightarrow v_B\sin\phi=v_A\sin\theta\qquad ...........(2)"

Squaring (1) and (2) and adding them together,

"v^2_B\cos^2\phi+v^2_B\sin^2\phi =(u_A-v_A\cos\theta )^2+v^2_A\sin^2\theta\\
\Rightarrow v^2_B(\cos^2\phi+\sin^2\phi)=(u^2_A-2u_Av_A \cos\theta+v^2_A\cos^2\theta)+v^2_A\sin^2\theta\\
\Rightarrow v^2_B(\cos^2\phi+\sin^2\phi)=u^2_A-2u_Av_A \cos\theta+v^2_A(\cos^2\theta+\sin^2\theta)\\
\Rightarrow v^2_B=u^2_A-2u_Av_A \cos\theta+v^2_A"

where we have used the property "\sin^2x+cos^2x=1" .

"\therefore v_B=\sqrt{ u^2_A-2u_Av_A \cos\theta+v^2_A}" ..............(3)

Given "u_A=3.00\times 10^2\ m/s"

"v_A=1.50\times 10^2\ m/s"

"\theta=60\degree"

Substituting these values in equation (3), we get

"v_B=\sqrt{ (3.00\times 10^2)^2-2\times 3.00\times 10^2\times 1.50\times 10^2\times \cos60\degree+(1.50\times 10^2)^2}"

"\Rightarrow v_B=2.56\times 10^2\ m/s"

Thus, the speed of the proton B after collision is "2.56\times 10^2\ m/s" .

Kinetic energy:

Total kinetic energy before collision = "\frac{1}{2}mu^2_A"

Total kinetic energy after collision "=\frac{1}{2}mv^2_A+\frac{1}{2}mv^2_B"

For the collision to be elastic, the total kinetic energy before collision must be equal to the total kinetic energy after collision i.e, for elastic collision

"\frac{1}{2}mu^2_A=\frac{1}{2}mv^2_A+\frac{1}{2}mv^2_B\\
\Rightarrow u^2_A=v^2_A+v^2_B"

If this condition holds, the collision is elastic. Otherwise, the collision is inelastic.

Let's check it now

"u^2_A=(3.00\times 10^2)^2=9.00\times 10^4"

"v^2_A+v^2_B=(1.50\times 10^2)^2+(2.56\times 10^2)^2=8.80\times 10^4"

Thus we can see that "u^2_A>(v^2_A+v^2_B)" which implies loss of energy and hence the collision is inelastic.


Answer: (i) The velocity of proton B after collision is "2.56\times 10^2\ m/s"

(ii) The collision is not elastic.